A Liquid mixture containing 1 mole benzene and 1 mole of toulene . The vapour pressure of pure benzene and toulene is 103 and 32.05 mm of hg.Then find out the pressure ,the composition of liquid ,the composition of vapour when 1 mole of mixture is vapourized?

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
mole fraction of benzene=1/2 and that of toluene=1/2
we have partial pressure of benzene=1/2*103 mm=51.5 mm
and partial pressure of toluene=.5*32.05=16.025 mm
therefore total pressure=sum of the partial pressure=51.5+16=67.525 mm
composition in the vapour phase=partial pressure/total=.76
mole fraction of toluene=16/67.525=.23
actually i have considered as liquid mixture whose mole of benzne and toluene are 1 mole respectively.
I f not to consider then total pressure data to be provided.